# Comparing invariant-mass spectroscopy of 8B with ab initio predictions

### v1

@@ IV Theory
The ab initio SA-NCSM is based on the no-core shell-model concept, but formulated with an -coupled or -coupled symmetry-adapted basis (see, e.g., Refs. [LauneyDD16, DytrychLDRWRBB20] and references therein). This enables physically relevant selections of the SA-NCSM model spaces, which in turn allows the nucleons to occupy much larger spaces. This is essential to accommodate large deformation, clustering and coupling to the continuum. In the SA-NCSM with continuum [LauneyMD_ARNPS21], the interior wave function is obtained from high-quality SA-NCSM calculations, whereas the asymptotics in the exterior can be straightforwardly recovered by using the exact Coulomb wavefunctions at large distances through, e.g., a microscopic R-matrix technique (currently implemented for partitionings of two clusters). In this way, the SA-NCSM framework can be utilized to provide structure observables, including transition strengths for Mg isotopes [Ruotsalainen19, launeyPPNP2026] and beta decays [Sargsyan_A8], and in addition, reaction observables, such as partial widths and asymptotic normalization coefficients [DreyfussLESBDD20, Sargsyan_A8, SargsyanLSMD2023], response functions and sum rules [BurrowsEWLMNP19, Baker:2024], as well as alpha capture and knock-out reaction cross sections [DreyfussLESBDD20, Sargsyan:2025].
@@
Similar to the conventional NCSM [NavratilVB00, NavratilVB00b], the SA-NCSM employs the single-particle basis of a harmonic oscillator (HO) potential, where major shells are separated by the parameter . The many-body model space is capped at an cutoff, which defines the maximum total number of oscillator quanta above the valence-shell configuration. As approaches infinity, the results converge toward the exact values and become independent of the HO parameter . The SA-NCSM utilizes a non-relativistic nuclear Hamiltonian, which includes translationally invariant inter-nucleon interaction alongside the Coulomb interaction between protons. SA-NCSM computes eigenvalues and eigenvectors of the nuclear Hamiltonian, which are subsequently used to evaluate nuclear observables. Since calculations are performed in laboratory coordinates, spurious center-of-mass excitations are removed from the low-lying spectrum via a Lawson term [Lawson74, DytrychMLDVCLCS11]. Within a given complete model space, the SA-NCSM results are equivalent to those obtained with the NCSM for the same interaction.
@@
The many-nucleon basis states of the SA-NCSM provide important information, namely, the total intrinsic spin , total orbital angular momentum , both coupled to , along with the intrinsic deformation. The deformation is informed by a pair of numbers , where and for a total of HO quanta distributed among the , , and directions (capped by ). The quantum numbers can be understood in the approximate mapping of the microscopic SU(3) basis states to a single rigid-rotor ellipsoid: The case where , or equivalently , describes a spherical configuration. A larger compared to (i.e., with ) indicates prolate quadrupole deformation; conversely, indicates oblate quadrupole deformation. Notably, a closed-shell configuration corresponds to , which is spherical (or no deformation).
@@
For the SA-NCSM calculations in this work, we adopt the NNLOopt chiral potential [Ekstrom:2013]. We perform calculations for 20 and 25 MeV, and for model spaces up to . For each of the values we perform extrapolations of the energies to obtain results by using a three-parameter exponential formula, similar to Ref. [MarisVS09]: . The extrapolated energies and their uncertainties are determined from across the range.
@@
The predicted level schemes of the two mirrors are quite similar with no large Thomas-Ehrman shifts which could be produced by significant double occupancy of the proton second single-particle orbital. The largest predicted shift of 0.6 MeV is found for the 0 state (considering centroid energies only). State energies from the Greens Function Monte Carlo (GFMC) approach [Piarulli:Priv] using the AV18+IL7 phenomenological interaction [Wiringa:1995, Pieper:2008] and the NV2+3-Ia chiral interaction [Piarulli:2018] are also provided in Table 3 for the 1+ states. The energies from these two calculations are similar to the SA-NCSM values although the agreement is notably better with the AV18+IL7 interaction whose levels are shown as the blue lines in Fig. 7. The energies of the low-lying excited states from SA-NCSM are also similar to the recent calculations from the no-core shell model with continuum for 7Be+p [kravvaris2023ab].
Spectroscopic factors and partial widths can be calculated from the ab initio SA-NCSM wavefunctions following Ref. [SargsyanLSMD2023] for a single-nucleon fragment and Ref. [DreyfussLESBDD20] for a multi-nucleon fragment.
@@
The results are reported in Table 4 for MeV HO frequency. The estimates do not include cluster normalization, which was found negligible for Li in Ref. [SargsyanLSMD2023]. For the proton channel, the SFs and partial widths are calculated from SA-NCSM overlaps in =8 model spaces. The Li and 3HeLi channels use SA-NCSM wavefunctions of 8B calculated in =10 model spaces, along with overlaps calculated for leading and subleading shape contributions within 8B and the decay products. The leading shape contribution, e.g., for 8B, comes from the deformation listed in Table 3 and its vibrations (multiples of 2-shell 1-particle-1-hole excitations up through =10) (see Refs. [DreyfussLESBDD20, LauneyMD_ARNPS21] for details). We note that the Li width for the state and the 3HeLi width for the state are based on subleading contributions only. The widths for the prompt Li decay include -, -, and -waves, with and channels for the system. The total widths are reported as a sum over all 6Li and 3He channels, including the and sequential decays through the 7Be() states. To estimate the widths for these sequential decays, we use branching ratios of 4% () and 96% () for the 7Be() state, and 99% () and 1% () for the 7Be() state, calculated from the 7Be SA-NCSM wavefunctions in model spaces and using experimental threshold energies. These are consistent with results reported in the Gamow shell model [Fernandez:2023]. The 7Be() alpha branching ratio is underestimated compared to 4.6% reported in Ref. [McCray:1963], but this small difference appears to be inconsequential to the total width estimates.
@@ V Discussion
Based on the SA-NCSM predictions shown in Fig. 7, candidates levels for the two observed peaks at 8.4 MeV associated with 2+6Li and 3He+5Li+3He+ exit channels are the 1, 2 and the 3 states which have very similar predicted excitation energies and decay widths. In a companion paper [CharityShortArxiv], it is argued that the 2+6Li peak must be the 1 peak, with a total theoretical width [ MeV] consistent with the experimental value (see Tables 4 and 5). The +3He+ peak may be a second decay branch of the same state as the extracted excitation energies and widths are similar. If that is the case, then from the relative yields of the two peaks, the branching ratio to the +3He+ channel would be 1.2(4) times that of the 2 branch after correcting for detector efficiency. However, the SA-NCSM calculations do not predict any significant decay branch to either the +7Be(5/2) or 3He+5Li channels (Table 4) which, after sequential decay, will populate the +3He+ exit channel. Thus we consider the possibility that the 2 and the levels are the source of this yield.
Indeed, both of these states (2 and 3) have large spectroscopic factors and significant widths for proton decay to the 5/2 state in 7Be (Table 4) which sequentially decays to the observed +3He+ final state. They also have much larger 3He+5Li partial widths compared to that of the state. Theory predicts total decay widths of MeV for the 2 state and MeV for the 3 state which are consistent with the experimental result within the uncertainties of both experiment and theory.
Finally, a wide [=2.44(53) MeV] state at =10.08 MeV was needed to fit the 2+6Li invariant-mass spectra. This proposed state is not well resolved in the invariant-mass spectra of Fig. 1 and the excess yield above background could also be fit assuming more than one contributing level. It is interesting to explore whether a wide state is expected at this energy. The 2 and 3 levels are potential candidates but, as just discussed, they are expected to populate the +3He+ channel. The 3 state predicted at =10.8 MeV is a clear candidate. This state is mostly an rotation of the 1 state but with the addition of non negligible configurations. Predicted partial decay widths, listed in Table 4, are dominated by sequential 2 decay through the 5/2 state in 7Be which populates the observed exit channel. The predicted prompt 2 decay width is suppressed compared to the 1 state as now wave emission is required and the significant =2 component reduces the overlap with the ground state of 6Li. There is also a +7Be(5/2) branch which will give rise to a smaller yield in the +3He+ exit channel. The total predicted decay width of 1.7(2) MeV is consistent with the experimental value of 2.44(53) MeV within the uncertainties of both. Thus it is plausible to assign some, or all, of the excess yield at 10.08 MeV to the 3 state.
@@ VI CONCLUSION
The structure of 8B and the mirror states in 8Li have also been studied theoretically in an ab initio approach using the symmetry-adapter no-core shell model with continuum. Excitation energies and spectroscopic factors have been calculated for 8B states covering the experimental range of excitation energies with spin/parities consistent with proton -wave knockout from 9C projectiles. This information was used to associate all of the observed resonances to theoretical counterparts. All the new levels can be associated with positive parity states expected to be produced following -wave proton knockout from a 9C projectile. From the present data, there is no need to invoke any proton -wave knockout producing negative-parity states.
@@ Acknowledgements.
This work is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG02-87ER-40316, DE-SC0023532, and under the FRIB Theory Alliance award DE-SC0013617. This work is also supported in part by the National Nuclear Security Administration through the Center for Excellence in Nuclear Training and University Based Research (CENTAUR) under grant number DE-NA-0004150. This work benefited from high performance computational resources provided by LSU (www.hpc.lsu.edu), the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility at Lawrence Berkeley National Laboratory operated under Contract No. DE-AC02-05CH11231, as well as the Frontera computing project at the Texas Advanced Computing Center, made possible by National Science Foundation award OAC-1818253.Appendix A Background for the 2+6Li events
@@
A number of processes not associated with 8B decay can contribute to the detected yield of 2+6Li events. These other processes contribute to the background upon which the resonance peaks of interest sit. For the 9C beam, a proton knockout is required to produce 8B states. An example of a background process is if the knocked-out proton is emitted inside the angular acceptance of the detector and is detected, but one of the protons from 2 resonance decay is not. One can also consider the knockout of two protons forming a 7Be resonance that proton decays. If one of the two knocked-out protons and the resonance-decay proton are detected in coincidence with the 6Li residual, a background event is generated. Background is also generated if three protons are knocked out directly producing a 6Li residue and two of these knocked-out protons are detected. There are more ways of producing background 2+6Li events in the fragmentation of the 13O. This background is not totally uncorrelated as there may be contributions from the decay of 7Be resonances present. A procedure to estimate the background in the invariant-mass spectra was presented in Ref. [Charity:2023] involving weighted event mixing. Herein we follow the procedure suggested in that work for the 2 decay of 12O.
@@
The background was constructed by mixing a from one +6Li event with and 6Li fragments from a second such event. The added lone proton is like detecting a non-resonance proton produced in the knockout or fragmentation stage of the reaction. In order to use these manufactured data, one has to weight the mixed events to suppress the background distribution at the smaller invariant masses. This suppression can be due to Coulomb correlations present in the fragmentation process. The weighting depends on the relative energy between the and 6Li fragments from the different events. In Ref. [Charity:2023] it was proposed that the weighting function should be taken as the measured correlation function determined for +core or +core events as the Coulomb barriers are very similar to the +core case. Furthermore, the and clusters produced in the fragmentation of a proton-rich projectile are generally not produced via resonant decay, but rather are created directly in the fragmentation process and so their correlation function would be free of resonances. However in this case, +6Li and +6Li events can also be produced from 8,9Be resonances. Indeed, in the 13O data set, the correlation functions shown in Fig. 8, obtained by dividing the experimental distributions by their mixed counterparts, show hints of resonances. The presence of the resonances is most pronounced for the +6Li correlation function, see Fig. 8(a), but there are also suggestions of resonances in the +6Li events, see Fig. 8(b).
@@
In order to remove the contribution from these resonances, leaving just the smoother Coulomb correlations, both correlation functions were fit with two resonances, dotted-green curves in Fig. 8. The intrinsic lineshapes of these resonances were parameterized with the -matrix formalism. These resonances sit upon the desired smooth weighting function of relative energy , solid-blue curves in Fig. 8, parameterized by
@@
For the +6Li data, the fitted peaks correspond to 8Be excited states at =23.0(1) and 25.0(1) MeV with intrinsic widths of 0.5 and 0.62(6) MeV, respectively. The former is possibly a decay branch of the =22.98(1)-MeV (=0.23(5) MeV) state listed in the ENSDF data base [ENSDF].
@@
For the +6Li data, there are indications of peaks at excitations energies 18.75 and 20.7 MeV in 9Be. The fitted values of are similar, i.e. 1.05(4) and 0.86(4) MeV for the +6Li and +6Li data, respectively. These are similar to the values of =1.0(2) and 1.5(2) MeV obtained from fitting the +8B and +9B correlations functions in Ref. [Charity:2023], respectively, obtained from the same data set.
@@ Appendix B Background for the +3He+ events
Given the width of the 7Be(7/2-) state is not that large (=175 keV [ENSDF]), the +3He+ events gated on the presence of this state (gate ) are a good example of sequential decay. In constructing the background for these sequential events, we have mixed a proton from one gated +3He+ event with the identified 7Be(7/2-)He+ products from another such event. These mixed events are then weighted based on the -7Be relative energy using Eq. 1. The uncertainty analysis considers the same range of the parameter as used in the 2+6Li events.

### v2

@@ IV Theory
The ab initio SA-NCSM is based on the no-core shell-model concept, but formulated with an -coupled or -coupled symmetry-adapted basis (see, e.g., Refs. [33, 22] and references therein). This enables physically relevant selections of the SA-NCSM model spaces, which in turn allows the nucleons to occupy much larger spaces. This is essential to accommodate large deformation, clustering and coupling to the continuum. In the SA-NCSM with continuum [35], the interior wave function is obtained from high-quality SA-NCSM calculations, whereas the asymptotics in the exterior can be straightforwardly recovered by using the exact Coulomb wavefunctions at large distances through, e.g., a microscopic R-matrix technique (currently implemented for partitionings of two clusters). In this way, the SA-NCSM framework can be utilized to provide structure observables, including transition strengths for Mg isotopes [47, 36] and beta decays [48], and in addition, reaction observables, such as partial widths and asymptotic normalization coefficients [21, 48, 49], response functions and sum rules [10, 3], as well as alpha capture and knock-out reaction cross sections [21, 50].
@@
Similar to the conventional NCSM [42, 43], the SA-NCSM employs the single-particle basis of a harmonic oscillator (HO) potential, where major shells are separated by the parameter . The many-body model space is capped at an cutoff, which defines the maximum total number of oscillator quanta above the valence-shell configuration. As approaches infinity, the results converge toward the exact values and become independent of the HO parameter . The SA-NCSM utilizes a non-relativistic nuclear Hamiltonian, which includes translationally invariant inter-nucleon interaction alongside the Coulomb interaction between protons. SA-NCSM computes eigenvalues and eigenvectors of the nuclear Hamiltonian, which are subsequently used to evaluate nuclear observables. Since calculations are performed in laboratory coordinates, spurious center-of-mass excitations are removed from the low-lying spectrum via a Lawson term [29, 23]. Within a given complete model space, the SA-NCSM results are equivalent to those obtained with the NCSM for the same interaction.
@@
The many-nucleon basis states of the SA-NCSM provide important information, namely, the total intrinsic spin , total orbital angular momentum , both coupled to , along with the intrinsic deformation. The deformation is informed by a pair of numbers , where and for a total of HO quanta distributed among the , , and directions (capped by ). The quantum numbers can be understood in the approximate mapping of the microscopic SU(3) basis states to a single rigid-rotor ellipsoid: The case where , or equivalently , describes a spherical configuration. A larger compared with (i.e., with ) indicates prolate quadrupole deformation; conversely, indicates oblate quadrupole deformation. Notably, a closed-shell configuration corresponds to , which is spherical (or no deformation).
@@
For the SA-NCSM calculations in this work, we adopt the NNLOopt chiral potential [24]. We perform calculations for 20 and 25 MeV, and for model spaces up to . For each of the values we perform extrapolations of the energies to obtain results by using a three-parameter exponential formula, similar to Ref. [38]: . The extrapolated energies and their uncertainties are determined from across the range.
@@
The predicted level schemes of the two mirrors are quite similar with no large Thomas-Ehrman shifts which could be produced by significant double occupancy of the proton second single-particle orbital. The largest predicted shift of 0.6 MeV is found for the 0 state (considering centroid energies only). State energies from the Greens function Monte Carlo (GFMC) approach [45] using the AV18+IL7 phenomenological interaction [55, 46] and the NV2+3-Ia chiral interaction [44] are also provided in Table 3 for the 1+ states. The energies from these two calculations are similar to the SA-NCSM values although the agreement is notably better with the AV18+IL7 interaction whose levels are shown as the blue lines in Fig. 7. The energies of the low-lying excited states from SA-NCSM are also similar to the recent calculations from the no-core shell model with continuum for 7Be+p [32].
Spectroscopic factors and partial widths can be calculated from the ab initio SA-NCSM wavefunctions following Ref. [49] for a single-nucleon fragment and Ref. [21] for a multi-nucleon fragment.
@@
The results are reported in Table 4 for MeV HO frequency. The estimates do not include cluster normalization, which was found negligible for Li in Ref. [49]. For the proton channel, the SFs and partial widths are calculated from SA-NCSM overlaps in =8 model spaces. The Li and 3HeLi channels use SA-NCSM wavefunctions of 8B calculated in =10 model spaces, along with overlaps calculated for leading and subleading shape contributions within 8B and the decay products. The leading shape contribution, e.g., for 8B, comes from the deformation listed in Table 3 and its vibrations (multiples of two-shell one-particle-one-hole excitations up through =10; see Refs. [21, 35] for details). We note that the Li width for the state and the 3HeLi width for the state are based on subleading contributions only. The widths for the prompt Li decay include , , and waves, with and channels for the system. The total widths are reported as a sum over all 6Li and 3He channels, including the and sequential decays through the 7Be() states. To estimate the widths for these sequential decays, we use branching ratios of 4% () and 96% () for the 7Be() state, and 99% () and 1% () for the 7Be() state, calculated from the 7Be SA-NCSM wavefunctions in model spaces and using experimental threshold energies. These are consistent with results reported in the Gamow shell model [37]. The 7Be() alpha branching ratio is underestimated compared with 4.6% reported in Ref. [39], but this small difference appears to be inconsequential to the total width estimates.
@@ V Discussion
Based on the SA-NCSM predictions shown in Fig. 7, candidates levels for the two observed peaks at 8.4 MeV associated with 2+6Li and 3He+5Li+3He+ exit channels are the 1, 2 and the 3 states which have very similar predicted excitation energies and decay widths. In a companion paper [14], it is argued that the 2+6Li peak must be the 1 peak, with a total theoretical width [ MeV] consistent with the experimental value (see Tables 4 and 5). The +3He+ peak may be a second decay branch of the same state as the extracted excitation energies and widths are similar. If that is the case, then from the relative yields of the two peaks, the branching ratio to the +3He+ channel would be 1.2(4) times that of the 2 branch after correcting for detector efficiency. However, the SA-NCSM calculations do not predict any significant decay branch to either the +7Be(5/2) or 3He+5Li channels (Table 4) which, after sequential decay, will populate the +3He+ exit channel. Thus we consider the possibility that the 2 and the levels are the source of this yield.
Indeed, both of these states (2 and 3) have large spectroscopic factors and significant widths for proton decay to the 5/2 state in 7Be (Table 4) which sequentially decays to the observed +3He+ final state. They also have much larger 3He+5Li partial widths compared with that of the state. Theory predicts total decay widths of MeV for the 2 state and MeV for the 3 state which are consistent with the experimental result within the uncertainties of both experiment and theory.
Finally, a wide [=2.44(53) MeV] state at =10.08 MeV was needed to fit the 2+6Li invariant-mass spectra. This proposed state is not well resolved in the invariant-mass spectra of Fig. 1 and the excess yield above background could also be fit assuming more than one contributing level. It is interesting to explore whether a wide state is expected at this energy. The 2 and 3 levels are potential candidates but, as just discussed, they are expected to populate the +3He+ channel. The 3 state predicted at =10.8 MeV is a clear candidate. This state is mostly an rotation of the 1 state but with the addition of non negligible configurations. Predicted partial decay widths, listed in Table 4, are dominated by sequential 2 decay through the 5/2 state in 7Be which populates the observed exit channel. The predicted prompt 2 decay width is suppressed compared with the 1 state as now wave emission is required and the significant =2 component reduces the overlap with the ground state of 6Li. There is also a +7Be(5/2) branch which will give rise to a smaller yield in the +3He+ exit channel. The total predicted decay width of 1.7(2) MeV is consistent with the experimental value of 2.44(53) MeV within the uncertainties of both. Thus it is plausible to assign some, or all, of the excess yield at 10.08 MeV to the 3 state.
@@ VI CONCLUSION
The structure of 8B and the mirror states in 8Li have also been studied theoretically in an ab initio approach using the symmetry-adapter no-core shell model with continuum. Excitation energies and spectroscopic factors have been calculated for 8B states covering the experimental range of excitation energies with spin and parities consistent with proton -wave knockout from 9C projectiles. This information was used to associate all of the observed resonances to theoretical counterparts. All the new levels can be associated with positive parity states expected to be produced following -wave proton knockout from a 9C projectile. From the present data, there is no need to invoke any proton -wave knockout producing negative-parity states.
@@ Acknowledgements.
This work is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG02-87ER-40316 and No. DE-SC0023532, and under the FRIB Theory Alliance Award No. DE-SC0013617. This work is also supported in part by the National Nuclear Security Administration through the Center for Excellence in Nuclear Training and University Based Research (CENTAUR) under Grant No. DE-NA-0004150. This work benefited from high performance computational resources provided by LSU (www.hpc.lsu.edu), the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility at Lawrence Berkeley National Laboratory operated under Contract No. DE-AC02-05CH11231, as well as the Frontera computing project at the Texas Advanced Computing Center, made possible by U.S. National Science Foundation Award No. OAC-1818253.Appendix A Background for the 2+6Li events
@@
A number of processes not associated with 8B decay can contribute to the detected yield of 2+6Li events. These other processes contribute to the background upon which the resonance peaks of interest sit. For the 9C beam, a proton knockout is required to produce 8B states. An example of a background process is if the knocked-out proton is emitted inside the angular acceptance of the detector and is detected, but one of the protons from 2 resonance decay is not. One can also consider the knockout of two protons forming a 7Be resonance that proton decays. If one of the two knocked-out protons and the resonance-decay proton are detected in coincidence with the 6Li residual, a background event is generated. Background is also generated if three protons are knocked out directly producing a 6Li residue and two of these knocked-out protons are detected. There are more ways of producing background 2+6Li events in the fragmentation of the 13O. This background is not totally uncorrelated because there may be contributions from the decay of 7Be resonances present. A procedure to estimate the background in the invariant-mass spectra was presented in Ref. [17] involving weighted event mixing. Herein we follow the procedure suggested in that work for the 2 decay of 12O.
@@
The background was constructed by mixing a from one +6Li event with and 6Li fragments from a second such event. The added lone proton is like detecting a non-resonance proton produced in the knockout or fragmentation stage of the reaction. In order to use these manufactured data, one has to weight the mixed events to suppress the background distribution at the smaller invariant masses. This suppression can be due to Coulomb correlations present in the fragmentation process. The weighting depends on the relative energy between the and 6Li fragments from the different events. In Ref. [17] it was proposed that the weighting function should be taken as the measured correlation function determined for +core or +core events because the Coulomb barriers are very similar to the +core case. Furthermore, the and clusters produced in the fragmentation of a proton-rich projectile are generally not produced via resonant decay, but rather are created directly in the fragmentation process and so their correlation function would be free of resonances. However in this case, +6Li and +6Li events can also be produced from 8,9Be resonances. Indeed, in the 13O dataset, the correlation functions shown in Fig. 8, obtained by dividing the experimental distributions by their mixed counterparts, show hints of resonances. The presence of the resonances is most pronounced for the +6Li correlation function, see Fig. 8(a), but there are also suggestions of resonances in the +6Li events, see Fig. 8(b).
@@
In order to remove the contribution from these resonances, leaving just the smoother Coulomb correlations, both correlation functions were fit with two resonances, dotted-green curves in Fig. 8. The intrinsic lineshapes of these resonances were parametrized with the -matrix formalism. These resonances sit upon the desired smooth weighting function of relative energy , solid-blue curves in Fig. 8, parameterized by
@@
For the +6Li data, the fitted peaks correspond to 8Be excited states at =23.0(1) and 25.0(1) MeV with intrinsic widths of 0.5 and 0.62(6) MeV, respectively. The former is possibly a decay branch of the =22.98(1)-MeV [=0.23(5) MeV] state listed in the ENSDF data base [1].
@@
For the +6Li data, there are indications of peaks at excitations energies 18.75 and 20.7 MeV in 9Be. The fitted values of are similar, i.e. 1.05(4) and 0.86(4) MeV for the +6Li and +6Li data, respectively. These are similar to the values of =1.0(2) and 1.5(2) MeV obtained from fitting the +8B and +9B correlations functions in Ref. [17], respectively, obtained from the same dataset.
@@ Appendix B Background for the +3He+ events
Given the width of the 7Be(7/2-) state is not that large (=175 keV [1]), the +3He+ events gated on the presence of this state (gate ) are a good example of sequential decay. In constructing the background for these sequential events, we have mixed a proton from one gated +3He+ event with the identified 7Be(7/2-)He+ products from another such event. These mixed events are then weighted based on the -7Be relative energy using Eq. 1. The uncertainty analysis considers the same range of the parameter as used in the 2+6Li events.